# Implementing Errors on Errors: Bayesian vs Frequentist

### v1

@@ 2.1 Ordinary procedure derived from Bayesian inference
Theoretical ${\color[rgb]{1,0.2,0}\hat{\mu}}$, ${\color[rgb]{1,0.2,0}\boldsymbol{\hat{\sigma}}}$, ${\color[rgb]{1,0.2,0}\boldsymbol{\hat{r}}}$, ${\color[rgb]{1,0.2,0}\boldsymbol{\hat{\omega}}}$ ${\color[rgb]{1,0.2,0}\mu}$, ${\color[rgb]{1,0.2,0}\boldsymbol{\sigma}}$, ${\color[rgb]{1,0.2,0}\boldsymbol{r}}$, ${\color[rgb]{1,0.2,0}\boldsymbol{\omega}}$, ${\color[rgb]{1,0.2,0}\boldsymbol{v}}$
Experimental ${\color[rgb]{0,0.5,1}\boldsymbol{\hat{d}}}$, ${\color[rgb]{0,0.5,1}\boldsymbol{\hat{w}}}$, ${\color[rgb]{0,0.5,1}\boldsymbol{\hat{s}}}$ ${\color[rgb]{0,0.5,1}\boldsymbol{d}}$, ${\color[rgb]{0,0.5,1}\boldsymbol{w}}$, ${\color[rgb]{0,0.5,1}\boldsymbol{s}}$
@@
We also collectively write ${\color[rgb]{1,0.2,0}\boldsymbol{\sigma}}:=\Set{{\color[rgb]{1,0.2,0}\sigma_{i}}}_{i=1,\dots,N}$. Under the above assumption, the PDF for ${\color[rgb]{0,0.5,1}\boldsymbol{\hat{d}}}$ is given by
@@
Now we come back to the Bayesian approach. When one further assumes that the prior is uniform, $f_{\circ}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\right)=\text{const.}$., and that ${\color[rgb]{1,0.2,0}\boldsymbol{\sigma}}$ are exactly equal to the quoted stated uncertainties by the experimental collaborations,
@@
$f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0,0.5,1}\boldsymbol{d}},{\color[rgb]{1,0.2,0}\boldsymbol{\sigma}},f_{\circ}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\right)=\text{const.}\right)$. ) $\propto\prod_{i}\mathcal{N}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}d_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}\mu},{\color[rgb]{0,0.5,1}s_{i}^{2}}\right)\propto\mathcal{N}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0,0.5,1}\overline{d}},{\color[rgb]{0,0.5,1}\overline{s}^{2}}\right),$ (7)
@@ 2.2 Probabilistic modeling of skepticism
D’Agostini introduces the following ratios as alternative random variables to ${\color[rgb]{1,0.2,0}\boldsymbol{\hat{\sigma}}}$:
@@
follow the gamma distribution, whose PDF is given by222 We follow the standard convention in statistics, where the shape parameter is listed first, followed by the rate parameter. This contrasts with the convention used by the PDG [1], where the rate parameter appears first.
@@ 2.3 D’Agostini’s ratio
Figure 1: PDFs $g\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\omega}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)$ and $f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}r}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)$ are shown in the left and right panels, respectively. For simplicity, the $i$-dependence of ${\color[rgb]{1,0.2,0}\omega_{i}}$ and ${\color[rgb]{1,0.2,0}r_{i}}$ is omitted in this figure and caption. The nuisance parameters are set to D’Agostini’s recommended values ${\color[rgb]{0.03,1,0.25}k}=1.29$, ${\color[rgb]{0.03,1,0.25}\lambda}=0.589$ (solid), and to Cowan’s ${\color[rgb]{0.03,1,0.25}k}={\color[rgb]{0.03,1,0.25}\lambda}=1/4{\color[rgb]{0.03,1,0.25}\varepsilon^{2}}$ with ${\color[rgb]{0.03,1,0.25}\varepsilon}=0.2$ (dashed), $0.5$ (dotted), and $1$ (dot-dashed), corresponding to ${\color[rgb]{0.03,1,0.25}k}={\color[rgb]{0.03,1,0.25}\lambda}=6.25$, $1$, and $0.25$, respectively. The associated numbers of virtual repetitions are ${\color[rgb]{0.03,1,0.25}n}=13.5$, $3$, and $1.5$. Among these, the choice ${\color[rgb]{0.03,1,0.25}\varepsilon}=0.5$ corresponds to our version of D’Agostini’s criterion $\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}}\right]=\operatorname{E}^{2}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}}\right]$.
@@
The former condition follows from the assumption that the standard deviations on average agree with the stated uncertainties, while the latter provides a conservative estimate. The constraints (22) are satisfied by
@@ 2.4 Likelihood function
For the likelihood function $f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}\boldsymbol{d}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}\mu}\right)$ in the Bayesian inference (2), we assume that the results are independent from each other and further that ${\color[rgb]{0,0.5,1}\hat{d}_{i}}$ follows a normal distribution with standard deviation ${\color[rgb]{1,0.2,0}r_{i}}{\color[rgb]{0,0.5,1}s_{i}}$ ($={\color[rgb]{1,0.2,0}\sigma_{i}}$ in the previous notation):
@@
This PDF exhibits the kernel of a Student’s $t$ distribution, reflecting marginalization over a gamma-distributed variance.444 Integrating the normal likelihood in Eq. (29) over the gamma prior yields a Student’s $t$ distribution with $\nu=2{\color[rgb]{0.03,1,0.25}k}$. Equation (30) matches its kernel up to reparameterization, exposing the heavy tails from variance uncertainty.
When one further assumes a uniform prior distribution for ${\color[rgb]{1,0.2,0}\hat{\mu}}$, the resulting posterior takes the form
@@
[…]
@@ 3.1 Cowan’s sampling scheme
In the Bayesian approach, the “true value” ${\color[rgb]{1,0.2,0}\hat{\mu}}$ is treated as a theoretical random variable that follows a prior distribution $f_{\circ}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\right)$ and is updated to a posterior distribution $f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0,0.5,1}\boldsymbol{d}},{\color[rgb]{0,0.5,1}\boldsymbol{s}}\right)$ upon observing data. In contrast, the frequentist approach considers ${\color[rgb]{1,0.2,0}\mu}$ to be the fixed “true value”, while the observed variables fluctuate around it according to Eq. (5).
To handle the “errors on errors” within the frequentist framework, Cowan provides a way to treat each estimated variance ${\color[rgb]{0,0.5,1}\hat{w}_{i}}$ as an experimental random variable fluctuating around a so-to-say “true variance” ${\color[rgb]{1,0.2,0}v_{i}}$. To implement this idea, Cowan introduces ${\color[rgb]{0.03,1,0.25}n}$ (virtual) repetitions for each real experiment $i$, resulting in ${\color[rgb]{0.03,1,0.25}n}$ random variables for each: [^5] In principle, ${\color[rgb]{0.03,1,0.25}n}$ may vary across experiments, resulting in an $i$-dependent ${\color[rgb]{0.03,1,0.25}n_{i}}$. However, this dependence is typically neglected in practice [4, 5], corresponding to D’Agostini’s notion of democratic skepticism; see footnote 3. Here we also omit the $i$-dependence, as it can be trivially recovered if needed.
@@ 3.2 Errors-on-errors parameter
Cowan defines what is called the errors-on-errors parameter:666 We follow the notation ${\color[rgb]{0.03,1,0.25}\varepsilon}$ used in Ref. [6], rather than ${\color[rgb]{0.03,1,0.25}r}$ in Ref. [3], for the same quantity (not to be confused with D’Agostini’s ${\color[rgb]{1,0.2,0}r_{i}}$). Although this parameter may in principle depend on $i$, we omit the $i$-dependence while retaining the original name, “errors-on-errors,” which reflects its possible variability across experiments; see also footnotes 3 and 5.
@@
The spirit behind this definition can be explained as follows [3]: For ${\color[rgb]{0,0.5,1}\hat{s}_{i}^{2}}:={\color[rgb]{0,0.5,1}\hat{w}_{i}}$, the standard error propagation gives $\Delta{\color[rgb]{0,0.5,1}w_{i}}\simeq 2{\color[rgb]{0,0.5,1}s_{i}}\Delta{\color[rgb]{0,0.5,1}s_{i}}$. Cowan then expects the analogous relation between the standard deviations: $\sigma\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{w}_{i}}\right]\simeq 2\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]\sigma\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]$. If one were to approximate $\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{w}_{i}}\right]\sim\operatorname{E}^{2}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]$, then one would vaguely expect777 Note, however, that this naive treatment leads to $\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]\sim 0$, implying $\sigma\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]\sim 0$. This limitation can also be seen more explicitly: It suggests that ${1\over 2}{\sigma\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{w}_{i}}\right]\over\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{w}_{i}}\right]}=\frac{1}{2}\sqrt{\frac{\Braket{{\color[rgb]{0,0.5,1}\hat{s}_{i}^{4}}}}{\Braket{{\color[rgb]{0,0.5,1}\hat{s}_{i}^{2}}}^{2}}-1}$ should be close to ${\sigma\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]\over\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]}=\sqrt{\frac{\Braket{{\color[rgb]{0,0.5,1}\hat{s}_{i}^{2}}}}{\Braket{{\color[rgb]{0,0.5,1}\hat{s}_{i}}}^{2}}-1},$ which is unjustifiable.

### v2

@@ 2.1 Ordinary procedure derived from Bayesian inference
Theoretical $^{\color[rgb]{1,0.2,0}\hat{\mu}}$, $^{\color[rgb]{1,0.2,0}\boldsymbol{\hat{\sigma}}}$, $^{\color[rgb]{1,0.2,0}\boldsymbol{\hat{r}}}$, $^{\color[rgb]{1,0.2,0}\boldsymbol{\hat{\omega}}}$ $\color[rgb]{1,0.2,0}\mu$}, $\color[rgb]{1,0.2,0}\boldsymbol{\sigma}$}, $\color[rgb]{1,0.2,0}\boldsymbol{r}$}, $\color[rgb]{1,0.2,0}\boldsymbol{\omega}$}, $\color[rgb]{1,0.2,0}\boldsymbol{v}$}
Experimental $^{\color[rgb]{0,0.5,1}\boldsymbol{\hat{d}}}$, $^{\color[rgb]{0,0.5,1}\boldsymbol{\hat{w}}}$, $^{\color[rgb]{0,0.5,1}\boldsymbol{\hat{s}}}$ $\color[rgb]{0,0.5,1}\boldsymbol{d}$}, $\color[rgb]{0,0.5,1}\boldsymbol{w}$}, $\color[rgb]{0,0.5,1}\boldsymbol{s}$}
@@
We also collectively write ${\color[rgb]{1,0.2,0}\boldsymbol{\sigma}}:=\Set{{\color[rgb]{1,0.2,0}\sigma_{i}}}_{i=1,\dots,N}$. Under the above assumption, the PDF for $^{\color[rgb]{0,0.5,1}\boldsymbol{\hat{d}}}$ is given by
@@
Now we come back to the Bayesian approach. When one further assumes that the prior is uniform, $f_{\circ}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\right)=\text{const.}$,222 The uniform prior adopted here is improper and non-integrable, but this does not affect our subsequent analysis as long as the likelihood function (7) decays sufficiently fast. and that $\color[rgb]{1,0.2,0}\boldsymbol{\sigma}$} are exactly equal to the quoted stated uncertainties by the experimental collaborations,
@@
$f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0,0.5,1}\boldsymbol{d}},{\color[rgb]{1,0.2,0}\boldsymbol{\sigma}},f_{\circ}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\right)=\text{const.}\right)$ $\propto\prod_{i}\mathcal{N}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}d_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}\mu},{\color[rgb]{0,0.5,1}s_{i}^{2}}\right)\propto\mathcal{N}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0,0.5,1}\overline{d}},{\color[rgb]{0,0.5,1}\overline{s}^{2}}\right),$ (7)
@@ 2.2 Probabilistic modeling of skepticism
D’Agostini introduces the following ratios as alternative random variables to $^{\color[rgb]{1,0.2,0}\boldsymbol{\hat{\sigma}}}$:
@@
follow the gamma distribution, whose PDF is given by333 We follow the standard convention in statistics, where the shape parameter is listed first, followed by the rate parameter. This contrasts with the convention used by the PDG [1], where the rate parameter appears first.
@@ 2.3 D’Agostini’s ratio
Figure 1: PDFs $g\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\omega}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)$ and $f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}r}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)$ are shown in the left and right panels, respectively. For simplicity, the $i$-dependence of $\color[rgb]{1,0.2,0}\omega_{i}$} and $\color[rgb]{1,0.2,0}r_{i}$} is omitted in this figure and caption. The nuisance parameters are set to D’Agostini’s recommended values ${\color[rgb]{0.03,1,0.25}k}=1.29$, ${\color[rgb]{0.03,1,0.25}\lambda}=0.589$ (solid), and to Cowan’s ${\color[rgb]{0.03,1,0.25}k}={\color[rgb]{0.03,1,0.25}\lambda}=1/4{\color[rgb]{0.03,1,0.25}\varepsilon^{2}}$ with ${\color[rgb]{0.03,1,0.25}\varepsilon}=0.2$ (dashed), $0.5$ (dotted), and 11 (dot-dashed), corresponding to ${\color[rgb]{0.03,1,0.25}k}={\color[rgb]{0.03,1,0.25}\lambda}=6.25$, $1$, and $0.25$, respectively. The associated numbers of virtual repetitions are ${\color[rgb]{0.03,1,0.25}n}=13.5$, $3$, and $1.5$. Among these, the choice ${\color[rgb]{0.03,1,0.25}\varepsilon}=0.5$ corresponds to our version of D’Agostini’s criterion $\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}}\right]=\operatorname{E}^{2}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}}\right$].
@@
The former condition follows from the assumption that the standard deviations on average agree with the stated uncertainties, while the latter provides a conservative estimate.555 As discussed in Section 2.3.6 of Ref. [4], some published measurements may contain biases or underestimate uncertainties, potentially affecting D’Agostini’s assumption that quoted uncertainties are unbiased and conservatively estimated. The constraints (22) are satisfied by
@@ 2.4 Likelihood function
For the likelihood function $f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}\boldsymbol{d}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}\mu}\right)$ in the Bayesian inference (2), we assume that the results are independent from each other and further that $\color[rgb]{0,0.5,1}\hat{d}_{i}$} follows a normal distribution with standard deviation ${\color[rgb]{1,0.2,0}r_{i}}{\color[rgb]{0,0.5,1}s_{i}}$ ($\color[rgb]{1,0.2,0}\sigma_{i}$} in the previous notation):
@@
This PDF exhibits the kernel of a Student’s $t$ distribution, reflecting marginalization over a gamma-distributed variance.666 Integrating the normal likelihood in Eq. (29) over the gamma prior yields a Student’s $t$ distribution with $\nu=2{\color[rgb]{0.03,1,0.25}k}$. Equation (30) matches its kernel up to reparameterization, exposing the heavy tails from variance uncertainty.
When one further assumes a uniform prior distribution for $^{\color[rgb]{1,0.2,0}\hat{\mu}}$, the resulting posterior takes the form
@@
$f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0,0.5,1}\boldsymbol{d}},{\color[rgb]{0,0.5,1}\boldsymbol{s}}\right)$ $=\int\text{d}{\color[rgb]{0.03,1,0.25}k}\,\text{d}{\color[rgb]{0.03,1,0.25}\lambda}\,f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0,0.5,1}\boldsymbol{d}},{\color[rgb]{0,0.5,1}\boldsymbol{s}},{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right),$ (32)
@@ 3.1 Cowan’s sampling scheme
In the Bayesian approach, the “true value” $^{\color[rgb]{1,0.2,0}\hat{\mu}}$ is treated as a theoretical random variable that follows a prior distribution $f_{\circ}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\right)$ and is updated to a posterior distribution $f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0,0.5,1}\boldsymbol{d}},{\color[rgb]{0,0.5,1}\boldsymbol{s}}\right)$ upon observing data. In contrast, the frequentist approach considers $\color[rgb]{1,0.2,0}\mu$} to be the fixed “true value”, while the observed variables fluctuate around it according to Eq. (5).
To handle the “errors on errors” within the frequentist framework, Cowan provides a way to treat each estimated variance $\color[rgb]{0,0.5,1}\hat{w}_{i}$} as an experimental random variable fluctuating around a so-to-say “true variance” $\color[rgb]{1,0.2,0}v_{i}$}. To implement this idea, Cowan introduces $\color[rgb]{0.03,1,0.25}n$} (virtual) repetitions for each real experiment $i$, resulting in $\color[rgb]{0.03,1,0.25}n$} random variables for each: [^7] In principle, $\color[rgb]{0.03,1,0.25}n$} may vary across experiments, resulting in an $i$-dependent $\color[rgb]{0.03,1,0.25}n_{i}$}. However, this dependence is typically neglected in practice [5, 6], corresponding to D’Agostini’s notion of democratic skepticism; see footnote 4. Here we also omit the $i$-dependence, as it can be trivially recovered if needed.
@@ 3.2 Errors-on-errors parameter
Cowan defines what is called the errors-on-errors parameter:888 We follow the notation $\color[rgb]{0.03,1,0.25}\varepsilon$} used in Ref. [7], rather than $\color[rgb]{0.03,1,0.25}r$} in Ref. [3], for the same quantity (not to be confused with D’Agostini’s $\color[rgb]{1,0.2,0}r_{i}$}). Although this parameter may in principle depend on $i$, we omit the $i$-dependence while retaining the original name, “errors-on-errors,” which reflects its possible variability across experiments; see also footnotes 4 and 7.
@@
The spirit behind this definition can be explained as follows [3]: For ${\color[rgb]{0,0.5,1}\hat{s}_{i}^{2}}:={\color[rgb]{0,0.5,1}\hat{w}_{i}}$, the standard error propagation gives $\Delta{\color[rgb]{0,0.5,1}w_{i}}\simeq 2{\color[rgb]{0,0.5,1}s_{i}}\Delta{\color[rgb]{0,0.5,1}s_{i}}$. Cowan then expects the analogous relation between the standard deviations: $\sigma\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{w}_{i}}\right]\simeq 2\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]\sigma\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]$. If one were to approximate $\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{w}_{i}}\right]\sim\operatorname{E}^{2}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]$, then one would vaguely expect999 Note, however, that this naive treatment leads to $\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]\sim 0$, implying $\sigma\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]\sim 0$. This limitation can also be seen more explicitly: It suggests that ${1\over 2}{\sigma\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{w}_{i}}\right]\over\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{w}_{i}}\right]}=\frac{1}{2}\sqrt{\frac{\Braket{{\color[rgb]{0,0.5,1}\hat{s}_{i}^{4}}}}{\Braket{{\color[rgb]{0,0.5,1}\hat{s}_{i}^{2}}}^{2}}-1}$ should be close to ${\sigma\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]\over\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]}=\sqrt{\frac{\Braket{{\color[rgb]{0,0.5,1}\hat{s}_{i}^{2}}}}{\Braket{{\color[rgb]{0,0.5,1}\hat{s}_{i}}}^{2}}-1},$ which is unjustifiable.